# Sylvester domains and pro-$p$ groups

### v1

@@ 6.2. Graded modules over $\bm{\operatorname{Gr}S}$
We have shown in Corollary 4.12 that the ring $\operatorname{Gr}S$ is a skew power series ring over the graded ring $\operatorname{Gr}R$, the latter which can be identified with $\Lambda$. Hence, we shall see $\operatorname{Gr}S$ as isomorphic to $\Lambda[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]$ where $\widetilde{s}=\overline{s}=\overline{g-1}$. Moreover, since this isomorphism is compatible with the isomorphism $\operatorname{Gr}\mathbb{D}\simeq\mathcal{D}[\overline{t}][\overline{s};\overline{\sigma},\overline{\delta}]$ under the injective map $\operatorname{Gr}S\to\operatorname{Gr}\mathbb{D}$, the maps $\widetilde{\sigma}$ and $\widetilde{\delta}$ are restrictions of $\overline{\sigma}$ and $\overline{\delta}$ to $\operatorname{Gr}S$ respectively. Hence, there are unique extensions $\widetilde{\sigma},\widetilde{\delta}$ from $\Lambda$ to $\mathcal{D}$ that that make the following diagram commute:
{tikzcd}
@@
From this and the fact that $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]$ is an Ore domain ([Rob01, Thm. 1.2.9 and 2.1.15]), the same argument in [HLA22] gives us:
Proposition 6.7 (cf. [HLA22, Lem. 2.7 and 2.8]).
@@
(1) Since $\operatorname{Gr}S$ has global dimension at most $2$, every left or right module $M$ has a projective resolution of length at most $2$. Hence, the claim follows.
(2) Let $0\to P_{1}\to P_{0}\to\mathcal{D}\to 0$ be a projective resolution of the left $\Lambda$-module $\mathcal{D}$ over $\Lambda$. Applying the functor $\operatorname{Gr}S\otimes_{\Lambda}-$ to this sequence we obtain the sequence
@@
which is again exact by the freeness of $\operatorname{Gr}S$ over $\Lambda$ (by both sides!). Since each $\operatorname{Gr}S\otimes_{\Lambda}P_{i}$ is again projective, this is a projective resolution of $\operatorname{Gr}S\otimes_{\Lambda}\mathcal{D}$ over $\operatorname{Gr}S$. We know that $\operatorname{Gr}S\otimes_{\Lambda}\mathcal{D}$ is isomorphic to $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]$ as a left $\operatorname{Gr}S$-module by Lemma 6.6, so the claim for the left module $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]$ follows. The proof for $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]$ as a right $\operatorname{Gr}S$-module is analogous, exchanging the functor $\operatorname{Gr}S\otimes_{\Lambda}-$ for $-\otimes_{\Lambda}\operatorname{Gr}S$.
@@
where the first term vanishes by (2) and the third term vanishes by (1). Hence, $\operatorname{Ext}^{2}_{\operatorname{Gr}S}(A,B)=0$ for all $\operatorname{Gr}S$-modules $B$, and the claim follows from [Rot08, Prop. 8.6].
(4) Every finitely generated $\operatorname{Gr}S$-submodule of $\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}])$ is $\operatorname{Gr}S$-isomorphic to an $\operatorname{Gr}S$-submodule of $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]$ by taking the smallest common multiple of the denominators of a generating set. Hence, the claim follows from (3).
@@
$\simeq 0\text{, because }\mathcal{D}\text{ has dimension }0\,.$
@@
(***) $\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}])\otimes_{\operatorname{Gr}S}P_{*}\simeq\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}])\otimes_{\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]\otimes_{\operatorname{Gr}S}P_{*})\,.$
Since the Ore localization is flat by [Rob01, Prop. 2.1.16(ii)], the homology of the right hand side of (***6.2) is isomorphic to the homology of $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]\otimes_{\operatorname{Gr}S}P_{*}$. Hence, we get isomorphisms $\operatorname{Tor}_{i}^{\operatorname{Gr}S}(\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]),N)\simeq\operatorname{Tor}_{i}^{\operatorname{Gr}S}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}],N)$ for all $i\geq 0$. In particular, $\operatorname{Tor}_{1}^{GrS}(\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]),N)=0$ by (7).
@@ 6.3. $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain
Since every free-by-$\mathbb{Z}_{p}$ pro-$p$ group has cohomological dimension at most $2$, our ring $S=\mathbb{F}_{p}[\![G]\!]$ has “continuous” global dimension at most $2$ in the sense of A. Brumer ([Bru66, Sec. 3]). By the characterization of $P$ being a projective profinite $S$-module if and only if $\mathfrak{Tor}_{1}^{S}(\mathbb{F}_{p},P)=0$ given in [Bru66, Prop. 3.1], through dimension shifting we can improve the results of Proposition 6.1:
@@
Inducing a filtration on $K_{1}$ from $\widehat{S(V_{1})}$, it becomes again a separated, complete and filtered $S$-module with a strict inclusion map $K_{1}\to\widehat{S(V_{1})}$. Since $\widehat{S(V_{1})}\simeq S[\![(\bm{V_{1}}\cup\{\infty\},\infty)]\!]$ is countably based, then so is $K_{1}$, which we identify with a free profinite $S$-module $S[\![(X,*)]\!]$ on a countable profinite pointed set $(X,*)$. By counting how many elements of $X$ have a non-trivial image in each homogeneous component of $\operatorname{Gr}K_{1}$, one obtains a multiset of integers $V$ such that the induced map $S(V)\to K_{1}$ is injective and has a dense image. This implies that $\widehat{S(V)}\to K_{1}$ is an isomorphism, completing the proof. ∎
@@
We follow the ideas of [JZ20, Prop. 4.13]. We will prove the proposition for left modules only, since the proof for right modules is analogous. We can assume that $M\subseteq\mathbb{D}$, so that $M$ is a filtered, separated, bounded and complete $S$-module. Let us take
@@
{tikzcd}
where the vertical arrows are isomorphisms by [JZ20, Lem. 4.8]. Since we’ve shown the bottom row is exact, the same holds for the top row, showing by [JZ20][Prop. 4.4] that the sequence $\mathbb{D}\widehat{\otimes}_{S}\dagger$ is exact at $\widehat{\mathbb{D}(V_{i})}$ for $i\in\{1,2\}$. That the map $\operatorname{id}\widehat{\otimes}d_{0}$ is surjective is also clear, so it remains to show exactness at $\widehat{\mathbb{D}(V_{0})}$.
@@
Since $M$ is a finitely generated $S$-module, then $\mathbb{D}\widehat{\otimes}_{S}M$ is a finitely generated $\mathbb{D}$-module ([JZ20, Cor. 4.3]), so that
@@
On one hand, Shapiro’s isomorphism ([Rot08, Cor. 10.72]) gives us:
@@
$\dim_{\mathbb{F}_{p}}\ker(\widehat{\mathbb{F}_{p}(V_{n})}\to\widehat{\mathbb{F}_{p}(V_{n-1})})\leq\dim_{\mathcal{D}}\ker(\widehat{\mathcal{D}(V_{n})}\to\widehat{\mathcal{D}(V_{n-1})})$
@@
Let $M$ be a finitely generated left or right $S$-submodule of $\mathcal{Q}$. Then for any exact sequence $0\to I\to S^{d}\to M\to 0$ of left $S$-modules, $I$ is free of finite rank.
@@
Again, we will only prove the corollary for left $S$-submodules. Since $\mathbb{F}_{p}$ with trivial $G$-action is the only simple and pseudocompact $S$-module, by Proposition 6.11 and [Bru66, Cor. 3.2], we know that $M$ has projective dimension at most $1$ as an $S$-module. Hence, $I$ is projective by [Rot08, Prop. 8.6] and thus free by Kaplansky’s theorem on local rings. It’s rank as a free $S$-module is then given by
@@
Since $\mathcal{Q}$ is flat as a $\mathbb{D}$-module ([Rob01, Prop. 2.1.16]), if $P_{*}\to\mathcal{\mathcal{missing}}{Q}$ is a projective resolution of $\mathcal{Q}$ over $S$ we have the isomorphisms:
@@
We have shown in Proposition 6.11 that $M$ is of type $FP_{\infty}$. Since $N$ is bounded, the claim now follows from Theorem 6.5. ∎
@@ 7. The equality $\operatorname{rk}_{G}=\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}$
Proof.
For each $i\in\mathbb{Z}$ define following subsets of $G$:
@@
$w(\varphi(g-1))=\max\{w(\varphi(g)),w(-1))\}=v(\varphi(g))>1\,,$
@@
From Claim  7.2 it follows that the $G_{i}$ form a chain of normal open subgroups of $G$, for it is easily verified that each $G_{i}$ contains $1$ and is closed under products, inverses and conjugation. Since $w(q)=0$ if and only if $q=0$, we can also conclude that $\bigcap_{i\geq 0}G_{i}=\{1\}$, that is, the chain $G_{i}$ is residual in $G$. Hence, this chain can be used to compute the rank function $\operatorname{rk}_{G}$ of Example 2.5.
@@
Let $r=\sum_{i=1}^{n}q_{i}z^{i}$ and $r^{\prime}=\sum_{i=1}^{n}q_{i}^{\prime}z^{i}$ be two arbitrary elements in $\mathcal{Q}[z]$. We have $w^{\prime}(r)=0$ if and only if all of the $q_{i}$ have valuation 00, that is, if and only if $q_{i}=0$ for all $i$. Hence, $w^{\prime}(r)=0$ implies $r=0$. Suppose the minimal $w^{\prime}$ valuation on $r$ and $r^{\prime}$ is attained at the terms $q_{m}z^{m}$ and $q_{n}^{\prime}z^{n}$ with $m$ and $n$ minimal for this properties. Then, for all $0\leq i\,,j\leq n$, we have:
@@
Suppose now, by contradiction, that there exists $i\neq m$ and $j\neq n$ with $i+j=m+n$ such that $w(q_{i}q^{\prime}_{j})\geq w(q_{m}q^{\prime}_{n})$. Without loss of generality, we may suppose that $i<m$, so that $n<j$. By the minimality of $m$, we must have $w(q_{i})p^{-i}<w(q_{m})p^{-m}$, and by the definition of $n$ we have $w(q^{\prime}_{j})p^{-j}\leq w(q^{\prime}_{n})p^{-n}$. This implies:
@@
By virtue of Claim 7.3, we can suppose that $\mathcal{Q}$ contains a central element $z$ of valuation $p^{-1}$, for a division ring extension does not change the induced rank function on $\mathbb{F}_{p}[\![G]\!]$. Consider now $\mathcal{Q}_{0}$, the ring of elements with non-negative valuation. We note that $\varphi(\mathbb{F}_{p}[\![G]\!])\leq\mathcal{\mathcal{missing}}{Q}_{0}$ and the non-zero ideals of $\mathcal{Q}_{0}$ are all of the form $z^{i}\mathcal{Q}_{0}$ for $i\geq 0$. This implies that $\mathcal{Q}_{0}/z\mathcal{Q}_{0}$ is the only simple $\mathcal{Q}_{0}$-module, and the length of the $\mathcal{Q}_{0}$-modules $\mathcal{Q}_{0}/z^{i}\mathcal{Q}_{0}$ are precisely $i$.
@@
the length of the induced $\mathcal{Q}_{0}$-module. The proof henceforth will follow similar arguments to the ones in [JZ20, Sec. 5].
Let $\dim_{G/G_{i}}$ and $\varphi^{\#}\dim_{\mathcal{Q}}$ be the Sylvester module rank functions associated to $\operatorname{rk}_{G/G_{i}}$ and $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}$. Then, since it suffices to show that $\operatorname{rk}_{G}(A)=\lim_{i\to\infty}\operatorname{rk}_{G/G_{i}}(A)=n$, it only remains to show that
@@
It is immediate to verify that $\psi^{i}(q\otimes m)=qg^{i}\otimes g^{-i}m-q\otimes m$ and therefore $\psi^{p}=0$. Moreover, $L/\psi(L)\simeq\mathcal{Q}_{0}\otimes_{\mathbb{F}_{p}[\![V]\!]}M$, and by applying $\psi$ we find that this surjects onto $\psi^{i}(L)/\psi^{i+1}(L)$ for every $i$. Hence:
@@
Now we are ready to prove the theorem. Note that, by Lemma 2.4, to show that $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}\leq\operatorname{rk}_{G}$ it suffices to show that they coincide over square matrices invertible under $\varphi$. Indeed, any matrix $A\in\operatorname{Mat}_{n\times m}(\mathbb{F}_{p}[\![G]\!])$ contains a maximal submatrix $B\in\operatorname{Mat}_{k\times k}(\mathbb{F}_{p}[\![G]\!])$ such that

### v2

@@ 6.2. Graded modules over $\bm{\operatorname{Gr}S}$
We have shown in Corollary 4.12 that the ring $\operatorname{Gr}S$ is a skew polynomial ring over the graded ring $\operatorname{Gr}R$, the latter which can be identified with $\Lambda$. Hence, we shall see $\operatorname{Gr}S$ as isomorphic to $\Lambda[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}$] where $\widetilde{s}=\overline{s}=\overline{g-1}$. Moreover, since this isomorphism is compatible with the isomorphism $\operatorname{Gr}\mathbb{D}\simeq\mathcal{D}[\overline{t}][\overline{s};\overline{\sigma},\overline{\delta}]$ under the injective map $\operatorname{Gr}S\to\operatorname{Gr}\mathbb{D}$, the maps $\widetilde{\sigma}$ and $\widetilde{\delta}$ are restrictions of $\overline{\sigma}$ and $\overline{\delta}$ to $\operatorname{Gr}S$ respectively. Hence, there are unique extensions $\widetilde{\sigma},\widetilde{\delta}$ from $\Lambda$ to $\mathcal{D}$ that that make the following diagram commute:
$\mathcal{D}$$\mathcal{D}[t]$$\mathcal{D}$$\mathcal{D}[t]\,.$$\scriptstyle{\widetilde{\sigma},\widetilde{\delta}}$$\scriptstyle{\iota}$$\scriptstyle{\overline{\sigma},\overline{\delta}}$$\scriptstyle{\iota}$
@@
From this and the fact that $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}$] is an Ore domain ([43, Thm. 1.2.9 and 2.1.15]), the same argument in [17] gives us:
Proposition 6.7 (cf. [17, Lem. 2.7 and 2.8]).
@@
(1) Since the skew polynomial ring $\operatorname{Gr}S$ over the free algebra $\operatorname{Gr}R$ has global dimension at most $2$ by [43, Thm. 7.5.3] and [10, Cor. 2.5.2], every left or right module $M$ has a projective resolution of length at most $2$. Hence, the claim follows.
(2) Let $0\to P_{1}\to P_{0}\to\mathcal{D}\to 0$ be a projective resolution of the left $\Lambda$-module $\mathcal{D}$ over $\Lambda$, which exists because $\Lambda$ is a free algebra and hence has global dimension at most $1$. Applying the functor $\operatorname{Gr}S\otimes_{\Lambda}$- to this sequence we obtain the sequence
@@
which is again exact by the freeness of $\operatorname{Gr}S$ over $\Lambda$ (by both sides!). Since each $\operatorname{Gr}S\otimes_{\Lambda}P_{i}$ is again projective, this is a projective resolution of $\operatorname{Gr}S\otimes_{\Lambda}\mathcal{D}$ over $\operatorname{Gr}S$. We know that $\operatorname{Gr}S\otimes_{\Lambda}\mathcal{D}$ is isomorphic to $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}$] as a left $\operatorname{Gr}S$-module by Lemma 6.6, so the claim for the left module $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}$] follows. The proof for $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}$] as a right $\operatorname{Gr}S$-module is analogous, exchanging the functor $\operatorname{Gr}S\otimes_{\Lambda}$- for $\otimes_{\Lambda}\operatorname{Gr}S$.
@@
where the first term vanishes by (2) and the third term vanishes by (1). Hence, $\operatorname{Ext}^{2}_{\operatorname{Gr}S}(A,B)=0$ for all $\operatorname{Gr}S$-modules $B$, and the claim follows from [44, Prop. 8.6].
[…]
@@
[…]
@@
(*) $\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}])\otimes_{\operatorname{Gr}S}P_{*}\simeq\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}])\otimes_{\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]\otimes_{\operatorname{Gr}S}P_{*})\,.$
Since the Ore localization is flat by [43, Prop. 2.1.16(ii)], the homology of the right hand side of (*6.2) is isomorphic to the homology of $\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]\otimes_{\operatorname{Gr}S}P_{*}$. Hence, we get isomorphisms $\operatorname{Tor}_{i}^{\operatorname{Gr}S}(\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]),N)\simeq\operatorname{Tor}_{i}^{\operatorname{Gr}S}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}],N)$ for all $i\geq 0$. In particular, $\operatorname{Tor}_{1}^{GrS}(\operatorname{Ore}(\mathcal{D}[\widetilde{s};\widetilde{\sigma},\widetilde{\delta}]),N)=0$ by (7).
@@ 6.3. $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain
Since every free-by-$\mathbb{Z}_{p}$ pro-$p$ group has cohomological dimension at most $2$, our ring $S=\mathbb{F}_{p}[\![G]\!]$ has “continuous” global dimension at most $2$ in the sense of A. Brumer ([8, Sec. 3]). By the characterization of $P$ being a projective profinite $S$-module if and only if $\mathfrak{Tor}_{1}^{S}(\mathbb{F}_{p},P)=0$ given in [8, Prop. 3.1], through dimension shifting we can improve the results of Proposition 6.1:
@@
Inducing a filtration on $K_{1}$ from $\widehat{S(V_{1})}$, it becomes again a separated, complete and filtered $S$-module with a strict inclusion map $K_{1}\to\widehat{S(V_{1})}$. Since $\widehat{S(V_{1})}\simeq S[\![(\bm{V_{1}}\cup\{\infty\},\infty)]\!]$ is countably based, then so is $K_{1}$, which we identify with a free profinite $S$-module S[[(X,)]]S[\![(X,*)]\!] on a countable profinite pointed set (X,)(X,*). By counting how many elements of $X$ have a non-trivial image in each homogeneous component of $\operatorname{Gr}K_{1}$, one obtains a multiset of integers $V$ such that the induced map $S(V)\to K_{1}$ is injective and has a dense image. This implies that $\widehat{S(V)}\to K_{1}$ is an isomorphism, completing the proof. ∎
@@
We follow the ideas of [24, Prop. 4.13]. We will prove the proposition for left modules only, since the proof for right modules is analogous. We can assume that $M\subseteq\mathbb{D}$, so that $M$ is a filtered, separated, bounded and complete $S$-module. Let us take
@@
[…]
where the vertical arrows are isomorphisms by [24, Lem. 4.8]. Since we’ve shown the bottom row is exact, the same holds for the top row, showing by [24][Prop. 4.4] that the sequence $\mathbb{D}\widehat{\otimes}_{S}\dagger$ is exact at $\widehat{\mathbb{D}(V_{i})}$ for $i\in\{1,2\}$. That the map $\operatorname{id}\widehat{\otimes}d_{0}$ is surjective is also clear, so it remains to show exactness at $\widehat{\mathbb{D}(V_{0})}$.
@@
Since $M$ is a finitely generated $S$-module, then $\mathbb{D}\widehat{\otimes}_{S}M$ is a finitely generated $\mathbb{D}$-module ([24, Cor. 4.3]), so that
@@
On one hand, Shapiro’s isomorphism ([44, Cor. 10.72]) gives us:
@@
[…]
@@
Let $M$ be a finitely generated left or right $S$-submodule of $\mathcal{Q}$. Then for any exact sequence $0\to I\to S^{d}\to M\to 0$ of left $S$-modules, II is free of finite rank.
@@
Again, we will only prove the corollary for left $S$-submodules. Since $\mathbb{F}_{p}$ with trivial $G$-action is the only simple and pseudocompact $S$-module, by Proposition 6.11 and [8, Cor. 3.2], we know that $M$ has projective dimension at most $1$ as an $S$-module. Hence, $I$ is projective by [44, Prop. 8.6] and thus free by Kaplansky’s theorem on local rings. It’s rank as a free $S$-module is then given by
@@
Since $\mathcal{Q}$ is flat as a $\mathbb{D}$-module ([43, Prop. 2.1.16]), if $P_{*}\to\mathcal{\mathcal{}}{Q}$ is a projective resolution of $\mathcal{Q}$ over $S$ we have the isomorphisms:
@@
We have shown in Proposition 6.11 that $M$ is of type $FP_{\infty}$: given a presentation
$0\to K\to S^{n}\to M\to 0\,$,
with $n$ minimal we have that $K$ is finitely generated and thus profinite as $\operatorname{Tor}_{1}^{S}(M,\mathbb{F}_{p})=\operatorname{Tor}_{0}^{S}(K,\mathbb{F}_{p})$ is finite, an argument we can now repeat with $M$ replaced by $K$. Moreover, the vanishing of
$\mathfrak{Tor}_{1}^{S}(K,\mathbb{F}_{p})\simeq\operatorname{Tor}_{1}^{S}(K,\mathbb{F}_{p})\simeq\operatorname{Tor}_{2}^{S}(M,\mathbb{F}_{p})$
shows that it is projective by [8, Lem. 2.1(ii) and Prop. 3.1] and hence free. Since $S$ is bounded and $N$ is finitely generated, it is also bounded, and thus the claim now follows from Theorem 6.5. ∎
@@ 7. The equality $\operatorname{rk}_{G}=\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}$
We split again the proof into several lemmas. For each $i\in\mathbb{Z}$ define following subsets of $G$:
@@
$w(\varphi(g-1))=\max\{w(\varphi(g)),w(-1))\}=w(\varphi(g))>1\,$,
@@
By the continuity of $w$ it follows that the $G_{i}$ form a chain of normal open subgroups of $G$, for it is easily verified that each $G_{i}$ contains $1$ and is closed under products, inverses and conjugation. Since $w(q)=0$ if and only if $q=0$, we can also conclude that $\bigcap_{i\geq 0}G_{i}=\{1\}$, that is, the chain $G_{i}$ is residual in $G$. Hence, this chain can be used to compute the rank function $\operatorname{rk}_{G}$ of Example 2.5.
@@
Let $r=\sum_{i=1}^{n}q_{i}z^{i}$ and $r^{\prime}=\sum_{i=1}^{n}q_{i}^{\prime}z^{i}$ be two arbitrary elements in $\mathcal{Q}[z]$. We have $w^{\prime}(r)=0$ if and only if all of the $q_{i}$ have valuation 0, that is, if and only if $q_{i}=0$ for all $i$. Hence, $w^{\prime}(r)=0$ implies $r=0$. Suppose the minimal $w^{\prime}$ valuation on $r$ and $r^{\prime}$ is attained at the terms $q_{m}z^{m}$ and $q_{n}^{\prime}z^{n}$ with $m$ and $n$ minimal for this properties. Then, for all $0\leq i\,,j\leq n$, we have:
@@
Suppose now, by contradiction, that there exists $i\neq m$ and $j\neq n$ with $i+j=m+n$ such that $w(q_{i}q^{\prime}_{j})\geq w(q_{m}q^{\prime}_{n})$. Without loss of generality, we may suppose that $i<m$, so that $n<j$. By the minimality of $m$ and $n$, we must have $w(q_{i})p^{-i}<w(q_{m})p^{-m}$ and $w(q^{\prime}_{j})p^{-j}\leq w(q^{\prime}_{n})p^{-n}$. This implies:
@@
By virtue of Lemma 7.3, we can suppose that $\mathcal{Q}$ contains a central element $z$ of valuation $p^{-1}$, for a division ring extension does not change the induced rank function on $\mathbb{F}_{p}[\![G]\!]$. Consider now $\mathcal{Q}_{0}$, the ring of elements with non-negative valuation. We note that $\varphi(\mathbb{F}_{p}[\![G]\!])\leq\mathcal{\mathcal{}}{Q}_{0}$ and the non-zero left or right ideals of $\mathcal{Q}_{0}$ are the two-sided ideals $z^{i}\mathcal{Q}_{0}$ for $i\geq 0$: if $z^{i}u=uz^{i}\in\mathcal{Q}_{0}$ with $w(u)=1$, then $u^{-1}\in\mathcal{Q}_{0}$ and hence the ideals generated by $z^{i}u$ and $z^{i}$ coincide. This implies that $\mathcal{Q}_{0}/z\mathcal{Q}_{0}$ is the only simple $\mathcal{Q}_{0}$-module, and the length of the $\mathcal{Q}_{0}$-modules $\mathcal{Q}_{0}/z^{i}\mathcal{Q}_{0}$ are precisely $i$.
@@
the length of the induced $\mathcal{Q}_{0}$-module. The proof henceforth will follow similar arguments to the ones in [24, Sec. 5].
Let $\dim_{G/G_{i}}$ and $\varphi^{\#}\dim_{\mathcal{Q}}$ be the Sylvester module rank functions associated to $\operatorname{rk}_{G/G_{i}}$ and $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}$. Then, since it suffices to show that $\operatorname{rk}_{G}(A)=\lim_{i\to\infty}\operatorname{rk}_{G/G_{i}}(A)=n$, by the relation $\operatorname{rk}_{G}(A)=n-\dim_{G}(M)$ it only remains to show that
@@
It is immediate to verify that $\psi^{i}(q\otimes m)=\sum_{j=0}^{i}(-1)^{j}\binom{i}{j}qg^{i-j}\otimes g^{j-i}m$ and therefore $\psi^{p}=0$. Moreover, $L/\psi(L)\simeq\mathcal{Q}_{0}\otimes_{\mathbb{F}_{p}[\![V]\!]}M$, and by applying $\psi$ we find that this surjects onto $\psi^{i}(L)/\psi^{i+1}(L)$ for every $i$. Hence:
@@
Proof of Theorem 1.1.
Note that, by Lemma 2.4, to show that $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}\leq\operatorname{rk}_{G}$ it suffices to show that they coincide over square matrices invertible under $\varphi$. Indeed, any matrix $A\in\operatorname{Mat}_{n\times m}(\mathbb{F}_{p}[\![G]\!])$ contains a maximal submatrix $B\in\operatorname{Mat}_{k\times k}(\mathbb{F}_{p}[\![G]\!])$ such that